OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation. 3x + 2y = -21 x - 3y = 4

OpenStudy (anonymous):

a) 3(3y + 4) + 2y = -21 b) 3(-3y + 4) + 2y = -21 c) 3x + 2(3y + 4) = -21 d) 3x + 2(-3y + 4) = -21

OpenStudy (dls):

A

OpenStudy (anonymous):

is it really a? you said a for the last 2 answers

OpenStudy (dls):

yes